# for Web Agents

### v1

@@ 4.2 Main Results
Outperforming Tree-search and ReACT Methods:
As illustrated in Table 1, the R2D2 model achieves superior success rates across all tasks when compared to both the Tree-search and ReACT methods, demonstrating its effectiveness. For instance, in the CMS and Reddit tasks, R2D2 outperforms Tree-search by substantial margins (30.4% vs. 16.5% in CMS and 20.8% vs. 10.5% in Reddit). The improvement suggests that the integration of a systematic replay buffer and a reflective memory paradigm in R2D2 enhances the agent’s ability to navigate and learn from past interactions, avoiding repeated errors and optimizing the decision-making process based on agents’ experiences. This highlights the advantage of R2D2’s enhanced observability in the MDP, which enables a more comprehensive understanding of the web environment, leading to more accurate and efficient strategies.
Substantial Improvements in CMS and GitLab Domains:
The results in the CMS and GitLab domains are particularly notable. R2D2 achieved a 30.4% success rate in CMS and 28.0% in GitLab, considerably higher than other tested methods. These domains often require complex navigations with web interfaces, where R2D2’s capability to leverage past visited states and reflect on past actions proves especially beneficial.
Comparison with Complex Frameworks Like AR, LATS and AutoEval:
When compared to complex frameworks employing sophisticated reflection mechanisms (e.g., AR, AutoEval), R2D2 holds its own or exceeds performance, with a total success rate (SR) of 27.5%. While AR and AutoEval offer robust reflection capabilities, R2D2’s integrated approach to first remembering and then reflecting allows it to preemptively correct paths and further refine strategies. The success can be attributed to the method’s dual-paradigm system. We show more analysis in Section 4.3 and Section 5.
@@ 4.3 Ablation Study
Figure 4: Performance comparison with different reflection-based methods. R2D2 achieved marked increase at the first episode. Our manual inspection in Figure 6 shows 75% of the initial increase is attributed to fixing the navigation failures.
@@
Ablating Failed Trajectories. To elucidate the learning dynamics of R2D2, we conduct a study to isolate the impact of failed trajectories. During this study, only successful trajectories are provided as in-context demonstrations at inference time, thereby restricting R2D2 to learning exclusively from positive examples. This variant falls 7.5% from the full implementation of R2D2 to 20.5%. This also reveals a critical limitation: the number of positive examples is insufficient to provide robust navigation and reflection to the agent during inference. Consequently, if no relevant successful trajectory is identified at retrieval time. These findings substantiate the hypothesis that failed trajectories, despite not directly addressing user queries, are instrumental in enriching R2D2’s strategic repertoire, and R2D2 extend beyond the mere memorization of positive examples.
@@ 4.4 Error Analysis
As shown in Figure 6, we manually inspect the trajectories of the same 60 queries executed by the vanilla ReACT agent and R2D2 agent. About 60% of the vanilla ReACT agent trajectories stall at the navigation stage, so there is not even an opportunity to fail in execution. In contrast, the R2D2 agent substantially reduces navigation failures, reliably guiding itself toward the right content and thereby reaching a point where it is possible to fail in execution more frequently. As a result, R2D2 achieves a higher pass rate overall. We further annotate and discuss erroneous trajectories in Appendix A.
Figure 6: Error analysis of vanilla ReACT agents’ and R2D2’s trajectories. “E” indicates execution failures, “N” indicates navigation failures, and “S” indicates success. R2D2 substantially reduces navigation failures, achieving higher success rate.
5 Qualitative Study
We present a case study to conduct a qualitative evaluation of the R2D2 framework, as detailed in Appendix C. Our framework uses precise navigation and advanced reflective capabilities, enabling effective interaction with and extraction of information from intricate web environments. In contrast, the baseline struggles with navigation and data analysis, resulting in incomplete and inaccurate outcomes. These findings emphasize the importance of meticulous navigation strategies and reflective processing for successful information retrieval in sophisticated web interfaces.
@@ 6 Conclusion
The R2D2 framework significantly enhances web agents’ capabilities by integrating Remember and Reflect paradigms, enabling more effective navigation and interaction in complex web environments. This approach leads to measurable improvements in performance, reducing errors and increasing task completion rates. Through structured memory use and reflective learning, R2D2 not only outperforms existing models but also offers a scalable solution adaptable to various domains. Future work could extend its application, further optimizing agent functionality across broader scenarios.
@@
Our experiments were exclusively conducted in English. This limitation restricts our understanding of the model’s efficacy across different linguistic contexts, potentially overlooking cultural and language-specific nuances that could affect the agent’s performance in non-English web environments.
@@ Appendix B Prompt Details
Table 2: Comparison of action trajectories for resolving user queries. This table illustrates how the GPT-4o model fails to navigate complex web interfaces, whereas the proposed R2D2 framework successfully manages both intermediate steps and final actions to accurately address the same queries. Please refer to Section 5 for more detailed analysis.
@@
Figure 8: Trajectory evalution prompt for Update Operation. The LLM compares two trajectory and determines which one is better for addressing the user query.
@@ Appendix C Qualitative Analysis
In instances where navigation is executed correctly but is insufficient to solve the task, the reflection module of R2D2 plays a crucial role. A clear example of this is the task to list the top 2 search terms in the store. While the gpt-4o trajectory navigates correctly to the ’Search Terms’ section, it does not delve deeper into analyzing or sorting the data, resulting in incomplete and inaccurate information retrieval. Conversely, R2D2 not only accesses the correct section but also actively manipulates the data display by sorting the search terms according to their hits, thereby precisely identifying and articulating the top search terms. This demonstrates the power of R2D2’s reflective capabilities.

### v3

@@ 4.2 Main Results
Overall Performance. As shown in Table 1, our R2D2 model consistently achieves higher success rates than both Tree-Search and ReACT across all tasks. For example, on the CMS and Reddit tasks, R2D2 outperforms Tree-Search by substantial margins. These gains demonstrate the effectiveness of combining a systematic replay buffer with a reflective memory paradigm. By leveraging past interactions, R2D2 avoids repeated mistakes, leading to more accurate and efficient decisions.
Substantial Improvements in Complex Domains. The results in the CMS and GitLab domains are particularly notable. R2D2 achieved a 30% success rate in CMS and 28% in GitLab, considerably higher than other tested methods. These domains often require complex navigations with web interfaces, where R2D2’s capability to leverage past visited states and reflect on past actions proves especially beneficial.
Comparison with Reflection-based Frameworks. When compared to complex frameworks employing sophisticated reflection mechanisms (e.g., AR, AutoEval), R2D2 holds its own or exceeds performance, with a total success rate (SR) of 27.3%. While AR and AutoEval offer robust reflection capabilities, R2D2’s integrated approach to first remembering and then reflecting allows it to preemptively correct paths and further refine strategies. The success can be attributed to the method’s dual-paradigm system. We show more analysis in Section 4.3.
@@ 4.3 Ablation Study
Figure 4: Performance comparison with different reflection-based methods. R2D2 achieves marked increase at the first episode. Our manual inspection in Figure 6 shows 75% of the initial increase is attributed to fixing the navigation failures.
@@
Method Accuracy Steps
Tree-Search Koh et al. (2024) 19.2% 33.8
AutoEval Pan et al. (2024) 20.2% 29.2
R2D2 27.3% 13.1
Table 2: Comparison of task accuracy and number of actions required. R2D2 reduces the number of online steps while maintaining a higher success rate.
Ablating Failed Trajectories. To elucidate the learning dynamics of R2D2, we conduct a study to isolate the impact of failed trajectories. During this study, only successful trajectories are provided as in-context demonstrations at inference time, thereby restricting R2D2 to learning exclusively from positive examples. This variant falls 6.8% from the full implementation of R2D2 to 20.5%. This also reveals a critical limitation: the number of positive examples is insufficient to provide robust navigation and reflection to the agent during inference. Consequently, if no relevant successful trajectory is identified at retrieval time. These findings substantiate the hypothesis that failed trajectories, despite not directly addressing user queries, are instrumental in enriching R2D2’s strategic repertoire, and R2D2 extends beyond the mere memorization of positive examples.
4.4 Efficiency Analysis
Beyond improving task success rates, R2D2 also demonstrates significant efficiency gains by reducing the number of online steps required per task. As shown in Table 2, we compare the average steps taken to successfully address a task between R2D2 and open-sourced representative baselines. R2D2 completes tasks with fewer steps on average, achieving a higher success rate. [^6] Offline memory construction for R2D2 involves at most five actions per task, and the replay buffer creation is rule-based, making it lightweight. Because web-based tasks are often bottlenecked by interactions with the live environment rather than by language model queries, minimizing the number of online steps reduces latency and overall inference time. Consequently, R2D2 ’s ability to leverage a cached replay buffer avoids frequent back-and-forth rollouts, leading to improved efficiency alongside its superior performance.
@@ 4.5 Error Analysis
As shown in Figure 6, we manually inspect the trajectories of the same 60 queries executed by the vanilla ReACT agent and R2D2 agent. About 60% of the vanilla ReACT agent trajectories stall at the navigation stage, preventing the opportunity to fail in execution. In contrast, the R2D2 agent substantially reduces navigation failures, reliably guiding itself toward the right content and thereby reaching a point where it is possible to fail in execution more frequently. As a result, R2D2 achieves a higher pass rate overall. We further annotate and discuss erroneous trajectories in Appendix A. We also conduct a qualitative evaluation of the R2D2 framework, as detailed in Appendix C.
Figure 6: Error analysis of a vanilla ReACT agent’s and R2D2’s trajectories. “E” indicates execution failures, “N” indicates navigation failures, and “S” indicates success. R2D2 substantially reduces navigation failures, achieving a higher success rate.
@@ 5 Conclusion
The R2D2 framework significantly enhances web agents’ capabilities by integrating Remember and Reflect paradigms, enabling more effective navigation and interaction in complex web environments. This approach leads to measurable improvements in performance, reducing errors and increasing task completion rates. R2D2 not only outperforms existing models but also offers a scalable solution adaptable to various domains. Future work could extend its application, further optimizing agent functionality across broader scenarios.
Acknowledgments
Muhao Chen was supported by the DARPA FoundSci Grant HR00112490370, the NSF of the United States Grant ITE 2333736. This work is partially supported by DARPA award HR00112220046. Any opinions, findings, conclusions, or recommendations expressed here are those of the authors and do not necessarily reflect the view of our sponsors.
@@ Acknowledgments
Language Studied. Our experiments were exclusively conducted in English. This limitation restricts our understanding of the model’s efficacy across different linguistic contexts, potentially overlooking cultural and language-specific nuances that could affect the agent’s performance in non-English web environments.
Focus on a Single Benchmark. Our experiments are confined to the WebArena benchmark using GPT-4o, which may raise concerns about their broader applicability. However, WebArena spans a broad set of tasks, ranging from online shopping to social media interactions, and R2D2’s strong performance across these varied scenarios suggests that our cached-search approach and Remember/Reflect paradigms are not restricted to a single domain.
Resource Constraints. Each complete pass through WebArena incurs a cost of approximately \$200 in GPT-4o usage, making large-scale or multi-benchmark experimentation logistically challenging. That said, our approach does not inherently depend on GPT-4o. We anticipate that future research can replicate these methods using different LLM backends or other text-based environments, suggesting that our approach is not fundamentally limited in scope or generalizability.
@@ Appendix B Prompt Details
Table 3: Comparison of action trajectories for resolving user queries. This table illustrates how the GPT-4o model fails to navigate complex web interfaces, whereas the proposed R2D2 framework successfully manages both intermediate steps and final actions to accurately address the same queries.
@@
Figure 8: Trajectory evaluation prompt for Update Operation. The LLM compares two trajectory and determines which one is better for addressing the user query.
@@ Appendix C Qualitative Analysis
In instances where navigation is executed correctly but is insufficient to solve the task, the reflection module of R2D2 plays a crucial role. A clear example of this is the task to list the top two search terms in the store. While the gpt-4o trajectory navigates correctly to the ’Search Terms’ section, it does not delve deeper into analyzing or sorting the data, resulting in incomplete and inaccurate information retrieval. Conversely, R2D2 not only accesses the correct section but also actively manipulates the data display by sorting the search terms according to their hits, thereby precisely identifying and articulating the top search terms. This demonstrates the power of R2D2’s reflective capabilities.
Appendix D Further Efficiency Analysis of R2D2
The R2D2 framework’s efficiency and decision-making capabilities are significantly shaped by its two core components: the “Remember” and “Reflect” paradigms. This analysis will primarily focus on the “Remember” paradigm, which is responsible for dynamically reconstructing the web environment, effectively building a “map” that the agent uses for navigation. Its complexity is characterized by the replay buffer and the $A^{*}$ search algorithm.
Space Complexity of Replay Buffer. The replay buffer stores a graph where nodes represent observed web page states and edges represent transitions (actions). We would like to highlight that the replay buffer’s space complexity is $O(1)$, contending that its overall size is bounded by the number of website pages in a given domain. In practice, to ensure scalability even in potentially larger or more dynamic environments, we employ an eviction policy (e.g., removing least recently or least frequently visited nodes). This maintains a buffer of a fixed maximum size.
Time Complexity of $A*$ Search. The $A*$ search operates on the replay buffer graph. Its time complexity is generally $O(N\log N)$ when using an efficient priority queue, where $N$ is the number of nodes (states) expanded from the replay buffer graph during a specific search process. The search process, which relies on offline memory, requires fewer steps than other online search or reflection baselines as shown in Table 2. Because the primary time overhead is attributed to stepping within the web environment, we approximate the inference-time latency by counting the number of steps; And calls to the LLM contribute minimally to the overall time cost in our research implementation. Furthermore, in a production environment the LLM latency can be made nearly arbitrarily small, making the number of webpage steps even more significant to algorithm latency.